Three times the smallest of three consecutive even integers is two more than twice the largest. Find the integers.
Answer by checkley77(12844)

Let x, x+2 & x+4 be the three integers.
3x=(x+4)+2
3x=x+6
3x-x=6
2x=6
x=6/2
x=3 ans.
Proof:
3*3=(3+4)+2
9=7+2
9=9